Need help on finance!

Assume that it is now January 1, 2019. Wayne-Martin Electric Inc. (WME) has developed a solar panel capable of generating 200% more electricity than any other solar panel currently on the market. As a result, WME is expected to experience a 14% annual growth rate for the next 5 years. Other firms will have developed comparable technology by the end of 5 years, and WME's growth rate will slow to 6% per year indefinitely. Stockholders require a return of 11% on WME's stock. The most recent annual dividend (D0), which was paid yesterday, was $1.92 per share.

  1. Calculate WME's expected dividends for 2019, 2020, 2021, 2022, and 2023. Do not round intermediate calculations. Round your answers to the nearest cent.

    D2019 = $  

    D2020 = $  

    D2021 = $  

    D2022 = $  

    D2023 = $  

  2. Calculate the value of the stock today, . Proceed by finding the present value of the dividends expected at the end of 2019, 2020, 2021, 2022, and 2023 plus the present value of the stock price that should exist at the end of 2023. The year end 2023 stock price can be found by using the constant growth equation. Notice that to find the December 31, 2023, price, you must use the dividend expected in 2024, which is 6% greater than the 2023 dividend. Do not round intermediate calculations. Round your answer to the nearest cent. $   
  3. Calculate the expected dividend yield (D1/P0), capital gains yield, and total return (dividend yield plus capital gains yield) expected for 2019. (Assume that    and recognize that the capital gains yield is equal to the total return minus the dividend yield.) Do not round intermediate calculations. Round your answers to two decimal places.

    D1/P0 =   %

    Capital gains yield =   %

    Expected total return =   %

    Then calculate these same three yields for 2024. Do not round intermediate calculations. Round your answers to two decimal places.

    D6/P5 =   %

    Capital gains yield =   %

    Expected total return =   %

  4. How might an investor's tax situation affect his or her decision to purchase stocks of companies in the early stages of their lives, when they are growing rapidly, versus stocks of older, more mature firms? When does WME's stock become "mature" for purposes of this question?
    1. It is of no interest to investors whether they receive dividend income or capital gains income, since taxes on both types of income must be paid in the current year. The firm's stock is "mature" at the end of 2023.
    2. It is of no interest to investors whether they receive dividend income or capital gains income, since taxes on both types of income can be delayed until the stock is sold. The firm's stock is "mature" at the end of 2023.
    3. People in high-income tax brackets will be more inclined to purchase "growth" stocks to take the capital gains and thus delay the payment of taxes until a later date. The firm's stock is "mature" at the end of 2023.
    4. Some investors need cash dividends, while others would prefer growth. Investors must pay taxes each year on the capital gain during the year, while taxes on the dividends can be delayed until the stock is sold. The firm's stock is "mature" at the end of 2023.
    5. It is of no interest to investors whether they receive dividend income or capital gains income, since both types of income are always taxed at the same rate. The firm's stock is "mature" at the end of 2023.

    -Select one option
  5. Suppose your boss tells you she believes that WME's annual growth rate will be only 12% during the next 5 years and that the firm's long-run growth rate will be only 4%. Without doing any calculations, what general effect would these growth rate changes have on the price of WME's stock?
    1. Since the firm's supernormal and normal growth rates are lower, the dividends and, hence, the present value of the stock price will be lower. The total return from the stock will increase, and both the dividend yield and the capital gains yield will be greater than they were with the original growth rates.
    2. Since the firm's supernormal and normal growth rates are lower, the dividends and, hence, the present value of the stock price will be lower. The total return from the stock will decline, and the dividend yield and the capital gains yield will be the same.
    3. Since the firm's supernormal and normal growth rates are lower, the dividends and, hence, the present value of the stock price will be lower. The total return from the stock will remain the same, but the dividend yield will be larger and the capital gains yield will be smaller than they were with the original growth rates.
    4. Since the firm's supernormal and normal growth rates are lower, the dividends and, hence, the present value of the stock price will be lower. The total return from the stock will remain the same, but the dividend yield will be smaller and the capital gains yield will be larger than they were with the original growth rates.
    5. Since the firm's supernormal and normal growth rates are lower, the dividends and, hence, the present value of the stock price will be lower. The total return from the stock will decline, and both the dividend yield and the capital gains yield will be smaller than they were with the original growth rates.

    -Select one
  6. Suppose your boss also tells you that she regards WME as being quite risky and that she believes the required rate of return should be 17%, not 11%. Without doing any calculations, determine how the higher required rate of return would affect the price of the stock, the capital gains yield, and the dividend yield. Again, assume that the long-run growth rate is 4%.
    1. As the required return increases, the price of the stock goes up, and both the capital gains and dividend yields increase initially.
    2. As the required return increases, the price of the stock goes up, and both the capital gains and dividend yields decrease initially.
    3. As the required return increases, the price of the stock remains the same since both the capital gains and dividend yields remain constant.
    4. As the required return increases, the price of the stock goes down, but both the capital gains and dividend yields increase initially.
    5. As the required return increases, the price of the stock goes down, but both the capital gains and dividend yields decrease initially.

    -Select one
Answer #1

a. CALCULATION OF EXPECTED DIVIDEND:

D0

D2018

$1.92

D1=D0*1.14

D2019

$2.19

D2=D1*1.14

D2020

$2.50

D3=D2*1.14

D2021

$2.84

D4=D3*1.14

D2022

$3.24

D5=D4*1.14

D2023

$3.70

.b. Value of the stock today =Present Value of Future Cash Flow

D6=D2024=3.70*1.06=$3.92

Value of Stock in Year 5 =P5=D6/(R-g)

R= Required rate of return =11%=0.11

g= expected growth rate=6%=0.06

P5=3.92/(0.11-0.06)=$78.37

Present Value (PV) of Cash Flow =Cash Flow/((1+i)^N)

i=Required return =0.11

N=Year of Cah Flow

N

CF

CF/(1.11^N)

Year

Cash Flow

PV of Cash Flow

D1=D0*1.14

D2019

1

$2.19

1.971891892

D2=D1*1.14

D2020

2

$2.50

2.025186267

D3=D2*1.14

D2021

3

$2.84

2.079921031

D4=D3*1.14

D2022

4

$3.24

2.136135113

D5=D4*1.14

D2023

5

$3.70

2.193868495

P5=D6/(R-g)

P5

5

$78.37

46.50878058

SUM

56.91578338

Value of Stock today =$56.92

c. Dividend Yield in 2019 =D1/P0=2.19/56.92=0.0385=3.85%

Capital gain Yield =11-3.85=7.15%

Expected Total Return =11%

Yields for 2024

Dividend Yield =D6/P5=(3.70*1.06)/78.37=3.92/78.37=0.0500=5.00%

Capital gain Yield =11-5=6.00%

Expected Total Return =11%

.d. How might an investor's tax situation affect his or her decision to purchase stocks of companies in the early stages of their lives:

ANSWER:

III. People in high-income tax brackets will be more inclined to purchase "growth" stocks to take the capital gains and thus delay the payment of taxes until a later date. The firm's stock is "mature" at the end of 2023.

e. Expected future dividends will be lower. Consequently Future cash flows will be lower. Present Value of cash flows will be lower. Hence, Price of stock will be lower

ANSWER:

III.Since the firm's supernormal and normal growth rates are lower, the dividends and, hence, the present value of the stock price will be lower. The total return from the stock will remain the same, but the dividend yield will be larger and the capital gains yield will be smaller than they were with the original growth rates.

g. As required return increases, Present value of cash flow will be lower .Hence price will be lower.

ANSWER:

IV .As the required return increases, the price of the stock goes down, but both the capital gains and dividend yields increase initially.